OpenStudy (anonymous):

The local zoo has two water tanks for the elephant exhibit that are leaking One water tank contains 12 gal of water and is leaking at a constant rate of 3 gal/h. The second water tank contains 8 gal of water and is leaking at a constant rate of 5 gal/h. When will the two tanks have the same amount of water? Explain. Let x = the number of hours the tanks are filling and let y = the number of gallons in the tank.

OpenStudy (anonymous):

A. In –2 hours, because the solution to the system is (–2,18). B. They will never have the same amount of water because the solution to the system is (–2,18). It is not possible to have –2 gallons in the tanks. C. They will never have the same amount of water because the solution to the system is (–2,18). It is not possible to have time be –2 hours. D. In 2 hours, because the solution to the system is (2,18).

OpenStudy (anonymous):

@uri @PinkSapphire @aaggaa @P0sitr0n

OpenStudy (studybird123):

I would say B.

OpenStudy (anonymous):

Can you elaborate

OpenStudy (studybird123):

Oh wait its A.

OpenStudy (anonymous):

explain ..

OpenStudy (studybird123):

ok

OpenStudy (studybird123):

Well I have an older cousin they are in an online school and I was helping her out we both thought it was B. and it turns out it was A. and I had a friend who the thought the same thing and asked on here too this is the link http://openstudy.com/study#/updates/50696207e4b06e532421feca

OpenStudy (anonymous):

Well her choices are in a different order than mines so it would be D for me and I chose that the first time and it was wrong

jimthompson5910 (jim_thompson5910):

"The local zoo has two water tanks for the elephant exhibit that are leaking One water tank contains 12 gal of water and is leaking at a constant rate of 3 gal/h" so you start with 12 gallons and each hour, you're losing 3 gallons, which means hour1: 12 - 3*1 = 9 hour2: 12 - 3*2 = 6 hour3: 12 - 3*3 = 3 In general, at some unknown hour, t, you'll have 12 - 3t gallons left

jimthompson5910 (jim_thompson5910):

Similarly, when they say "The second water tank contains 8 gal of water and is leaking at a constant rate of 5 gal/h" this means the second tank will have 8 - 5t gallons of water after t hours

jimthompson5910 (jim_thompson5910):

They ask "When will the two tanks have the same amount of water?" so just set the two expressions equal to each other and solve for t so you need to solve the equation 12 - 3t = 8 - 5t

OpenStudy (anonymous):

t = 2

jimthompson5910 (jim_thompson5910):

you sure? you should check your answer

jimthompson5910 (jim_thompson5910):

you do so by plugging t = 2 back into the original equation

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

12 - 3t = 8 - 5t 12 - 3(2) = 8 - 5(2) ... replace EVERY copy of t with 2 12 - 6 = 8 - 10 6 = -2 <--- is this true?

OpenStudy (anonymous):

t = -2

jimthompson5910 (jim_thompson5910):

do you see why t = 2 doesn't work?

OpenStudy (anonymous):

yes it was a typo

jimthompson5910 (jim_thompson5910):

yes, t = -2

jimthompson5910 (jim_thompson5910):

what's the issue with saying t = -2?

OpenStudy (anonymous):

A ?

jimthompson5910 (jim_thompson5910):

does it make sense to have negative two hours?

OpenStudy (anonymous):

no lol

OpenStudy (anonymous):

so C ?

jimthompson5910 (jim_thompson5910):

correct, the two tanks will never have the same volumes at any given time

OpenStudy (anonymous):

Thank you so much !

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (studybird123):

Hey Shay Sorry I didn't answer correctly or give a good explanation.

OpenStudy (anonymous):

It's okay